Identical +4.78 µC charges are fixed to adjacent corners of a square. What charge (magnitude and algebraic sign) should be fixed to one of the empty corners, so that the total electric potential at the remaining empty corner is 0 V?
Answer #1

a = length of the side of square

r = length of the diagonal = sqrt(2) a

VA = electric potential at D due to charge at A = k Q/a

VB = electric potential at D due to charge at B = k Q/r

VC = electric potential at D due to charge at C = k q/a

Given that:

VA + VB + VC = 0

k Q/a + k Q/r + k q/a = 0

k Q/a + k Q/(sqrt(2) a) + k q/a = 0

Q + Q/(sqrt(2)) + q = 0

4.78 + 4.78/(sqrt(2)) + q = 0

q = - 8.16 C